OpenStudy (anonymous):

Graph the line with slope -1/3 passing through the point (4, -2) .

undeadknight26 (undeadknight26):

well u take ur first points and subtract 1 from the 4 and add 3 to the -2 i think! this should give u another point :D

OpenStudy (ybarrap):

Do you know the equation of a line?

OpenStudy (ybarrap):

Have you seen, \(y=mx+b\)?

OpenStudy (mathmale):

y = mx + b will do the trick, but in a roundabout way. Consider using the point-slope form of the equation of a straight line, y-y1 = m (x - x1), however. It's a lot faster for this type of problem. Just substitute the givens (slope and point (x1, y1)) into this equation, and then simplify the result to the form y = mx + b. Do you recall how to graph a line of the form y = mx + b? undeadknight presents an alternative way to find and graph a second point on the line in question. My first impulse was to use a closely related method: a) graph the given point, (4,-2) b) since the slope of the line is -1/3, representing "rise" over "run", count off 3 units in the (+) x direction from the given point and then count off 1 unit DOWNWARD in the (-) x direction, to locate the next point. Repeat this procedure. c) draw a line through these points.